### v2

# An Extension of the Euler-Maclaurin Summation Formula to Nearly Singular Functions

Abstract
A extension of the Euler-Maclaurin (E-M) formula to near-singular functions is presented. This extension is derived based on earlier generalized E-M formulas for singular functions. The new E-M formulas consists of two components: a “singular” component that is a continuous extension of the earlier singular E-M formulas, and a “jump” component associated with the discontinuity of the integral with respect to a parameter that controls near singularity. The singular component of the new E-M formulas is an asymptotic series whose coefficients depend on the Hurwitz zeta function or the digamma function. Numerical examples of near-singular quadrature based on the extended E-M formula are presented, where accuracies of machine precision are achieved insensitive to the strength of the near singularity and with a very small number of quadrature nodes.
1 Introduction
This paper considers simple and efficient numerical approximations of integrals of the form
$\int_{-a}^{a}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\,\mathrm{d}x$ (1)
where $a,c,d>0$ and $x_{s}\in(-a,a)$ are constants, and where $g\in C^{\infty}([-a,a])$. The regularity constraint on $g$ can be relaxed. When $d\approx 0$, the integrand is nearly singular at $x=x_{s}$, making the integral difficult to approximate using non-adaptive regular quadrature rules.
The consideration of (1) is motivated by the numerical challenge of evaluating integral operators of the form
$\int_{S}K(\tilde{\mathbf{x}},\mathbf{x})g(\mathbf{x})\mathrm{d}s_{\mathbf{x}}$ (2)
at a target point $\tilde{\mathbf{x}}$ that is at a distance $d\approx 0$ from the surface $S$, which leads to near-singularity of the type (1) in the kernel $K(\tilde{\mathbf{x}},\mathbf{x})$ [8]. The accurate and efficient evaluation of near-singular integrals is important in numerical simulations of physical systems that involve closely interacting components, such as viscous particulate flows or potential flows [3, 14].
The Trapezoidal rule with error corrections has proven to be effective for the discretization of integral operators such as (2) [20, 7, 2, 22]. The theoretical bases for these error corrections are the Euler-Maclaurin formula and its generalized forms. The classical Euler-Maclaurin formula [5, 10] provides an asymptotic error expansion for the trapezoidal quadrature rule applied to regular integrals. Starting from the 1960s, the Euler-Maclaurin formula has been generalized for integrals with integrable algebraic and logarithmic endpoint singularities by Navot [11, 12], for Hadamard finite-part integrals by Ninham and Lyness [13, 9], and finally for the $1/x$ type finite-part integrals by Sidi [18], who has also further extended the Euler-Maclaurin formula for polynomial compositions of algebraic and logarithmic endpoint singularities [17].
In this paper, we provide a new extension of the Euler-Maclaurin formula for nearly singular integrals of the form (1), which is derived from a combination of the above-mentioned singular Euler-Maclaurin formulas and the rational zeta series that draw connections between special functions. One unexpected consequence is an “off-mesh” Euler-Maclaurin formula for a hypersingular (i.e. finite-part) integral that is not seen in the literature before (see Corollary 10). It is also worth mentioning that the current work is closely connected to the work of Nitsche [14], in which the near-singular error corrections for the Trapezoidal rule are computed numerically (instead of using the analytic Euler-Maclaurin type expressions). While the results of the current paper provide more theoretical insights into the error analysis and makes the error correction easier to compute, the approach of Nitsche [14] can be more easily generalized to integrals in higher dimensions and has recently been applied to simulations of viscous fluid flows past ellipsoidal objects [15].
This paper is organized as follows. In Section 2, we first set up the notations and review the existing generalized Euler-Maclaurin formulas for singular integrals, which serves as the basis for approximating near-singular integrals. Then Section 3 presents the main result of this paper: a generalized Euler-Maclaurin formula for near-singular integrals of the form (1) and its analysis. Finally, numerical examples are given in Section 4, convergence of the new Euler-Maclaurin formula is demonstrated and compared with the classical regular Euler-Maclaurin formula. Further discussions and conclusions are in Section 5.
2 Preliminary: Euler-Maclaurin formulas for singular functions
We first set up the notations that will be used throughout the paper. Let $f$ be a function that is $C^{\infty}$ on $\mathbb{R}$ except for a possible singularity at $x=0$, where the $C^{\infty}$ regularity can be relaxed to $C^{p}$ for a sufficiently large $p$. Let $a>0$ and consider the following integrals
$I[f]$ $:=\int_{-a}^{a}f(x)\,\mathrm{d}x,$ $I_{L}[f]$ $:=\int_{-a}^{0}f(x)\,\mathrm{d}x,$ $I_{R}[f]$ $:=\int_{0}^{a}f(x)\,\mathrm{d}x,$ (3)
and their corresponding punctured Trapezoidal rules (where the possible singular point $x=0$ is omitted from the summation)
$L_{h}[f]$ $:=\sum_{k=-n}^{-1}f(kh)\,h+C_{h}^{-a},$ $R_{h}[f]$ $:=\sum_{k=1}^{n}f(kh)\,h+C_{h}^{a},$ $T_{h}[f]$ $:=L_{h}[f]+R_{h}[f],$ (4)
where $h=\frac{a}{n}$, and where $C_{h}^{-a}$ and $C_{h}^{a}$ are edge corrections at x=ax=-a and $x=a$, respectively, from the classical Euler-Maclaurin formula. For example, $C_{h}^{a}$ has the form
$C_{h}^{a}=-\frac{1}{2}f(a)h-\sum_{k=1}^{\lfloor p/2\rfloor}\frac{B_{2k}}{(2k)!}f^{(2k-1)}(a)h^{2k}$ (5)
where $B_{2k}$ are the Bernoulli numbers and where $p$ is sufficiently large such that the dominant quadrature error comes from the singularity of $f$ at $x=0$; alternatively, $C_{h}^{a}$ can also be approximated using Gregory interpolants [6]. An implementation of the Gregory interpolants is available in the Matlab code for the numerical examples in this paper, see Remark 5.
The error of the punctured Trapezoidal rule is
$E_{h}[f]:=I[f]-T_{h}[f].$ (6)
In addition, we also define the shifted Trapezoidal rules (such that $x=0$ is not a mesh point)
$L_{h,s}[f]$ $=\sum_{k=-n}^{-1}f(k\,h+s\,h)\,h+C_{h}^{-a+s}$ (7)
$R_{h,s}[f]$ $=\sum_{k=1}^{n}f(k\,h+s\,h)\,h+C_{h}^{a+s},$
$T_{h,s}[f]$ $=L_{h,s}[f]+R_{h,s}[f].$
where the shift is shs\,h for $|s|\leq\frac{1}{2}$; note that when $s=0$, (7) reduces to (4).
The following analytical tool will be critical for our analysis in the next section.
Theorem 1 (generalized Euler-Maclaurin formula).
Let $w\in\mathbb{C}$, then
$R_{h,s}\left[\frac{g(x)}{x^{w}}\right]-f.p.\,I_{R}\left[\frac{g(x)}{x^{w}}\right]\sim\sum_{k=0}^{\infty}\zeta_{h}(w-k,1+s)\frac{g^{(k)}(0)}{k!}h^{k+1-w}\quad\text{as }h\to 0,$ (8)
$L_{h,s}\left[\frac{g(x)}{x^{w}}\right]-f.p.\,I_{L}\left[\frac{g(x)}{x^{w}}\right]\sim\sum_{k=0}^{\infty}\zeta_{h}(w-k,1-s)(-1)^{k}\frac{g^{(k)}(0)}{k!}h^{k+1-w}\quad\text{as }h\to 0,$ (9)
where “f.p.f.p.” indicates that an integral is understood as a Hadamard finite-part integral, and where $\zeta_{h}(z,s)$ is a modified Hurwitz zeta function given by
$\zeta_{h}(z,s)=\begin{cases}\zeta(z,s),&z\neq 1\\ -\psi(s)-\log h,&z=1\end{cases}.$ (10)
where $\zeta(z,s)=\sum_{n=0}^{\infty}(n+s)^{-z}$ is the classical Hurwitz zeta function (which has an analytic extension for all $z\in\mathbb{C}\setminus\{1\}$), and where $\psi(s)={\Gamma^{\prime}(s)}/{\Gamma(s)}$ is the digamma function and $\Gamma(s)$ is the Euler’s gamma function.
For a proof of Theorem 1, see [11]. See also [4] for a summary of the properties of Hadamard finite-part integrals.
As a special case of Theorem 1, when $s=1$ we have the following corollary, whose proof can be found in [18].
Corollary 2.
$R_{h}\Big{[}\frac{g(x)}{x^{w}}\Big{]}-f.p.\,I_{R}\Big{[}\frac{g(x)}{x^{w}}\Big{]}\sim\sum_{k=0}^{\infty}\zeta_{h}(w-k)\frac{g^{(k)}(0)}{k!}h^{k+1-w}\quad\text{as }h\to 0,$ (11)
$L_{h}\Big{[}\frac{g(x)}{x^{w}}\Big{]}-f.p.\,I_{L}\Big{[}\frac{g(x)}{x^{w}}\Big{]}\sim\sum_{k=0}^{\infty}\zeta_{h}(w-k)(-1)^{k}\frac{g^{(k)}(0)}{k!}h^{k+1-w}\quad\text{as }h\to 0,$ (12)
$T_{h}\Big{[}\frac{g(x)}{x^{w}}\Big{]}-f.p.\,I\Big{[}\frac{g(x)}{x^{w}}\Big{]}\sim\sum_{k=0}^{\infty}2\zeta_{h}(w-2k)\frac{g^{(2k)}(0)}{(2k)!}h^{2k+1-w}\quad\text{as }h\to 0,$ (13)
where $\zeta_{h}(z)$ is a modified Riemann zeta function given by
$\zeta_{h}(z)=\begin{cases}\zeta(z),&z\neq 1\\ \gamma-\log h,&z=1\end{cases}.$ (14)
where $\zeta(z)$ is the classical Riemann zeta function, and where $\gamma=-\psi(1)$ is the Euler’s constant.
3 An Euler-Maclaurin formula for near-singular functions
The central focus of this section is the Trapezoidal quadrature approximation of the integral
$I\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}=\int_{-a}^{a}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\,\mathrm{d}x$ (15)
where $c,d>0$ are constants, $g(x)$ a smooth function compactly supported on [a,a][-a,a], and $x_{s}\in(-a,a)$ the location of near singularity. We will analyze the error of the punctured Trapezoidal rule $T_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}$; see Figure 1 for a schematic of the punctured Trapezoidal rule. Our analysis begins with the case $x_{s}=0$ such that the near-singularity is at a mesh point, namely $x=0$. Then we will consider the general case of $x_{s}\neq 0$. In both cases, we will show that the error $E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}$ consists of two components: a “singular error” which is a continuous extension of the generalized Euler-Maclaurin formulas from Section 2, and a “jump error” associated with the discontinuity of the value of the integral with respect to $d$.
Figure 1: Schematic of the punctured trapezoidal rule of mesh size $h$ for the near-singular integral (15), which can arise from an integral operator (2) with its target point $\tilde{\mathbf{x}}$ located at a distance $d$ from the surface $S$, whose orthogonal projection onto $S$ corresponds to $x_{s}=s\cdot h$ in the parameter space relative to the closest mesh point $x=0$. The open circle at $x=0$ indicates its omission in the summation of the punctured Trapezoidal rule.
3.1 Near-singular quadrature error analysis
Theorem 3 (Euler-Maclaurin for near-singular integral).
Given $g(x)\in C^{\infty}([a,a])$ and $c,d>0$, and using the notations (3-6), we have
$E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}x^{2}}\Big{]}$ $\sim-\frac{1}{c^{2}}\sum_{k=0}^{\infty}\frac{g^{(2k)}(0)}{(2k)!}h^{2k-1}(2\,z_{2k})+\frac{\pi}{cd}\sum_{k=0}^{\infty}\frac{g^{(2k)}(0)}{(2k)!}\left(-\frac{d^{2}}{c^{2}}\right)^{k}$ (16)
as $h\to 0$ and $d\to 0$, where the numbers $z_{k}$ satisfy the recurrence relation
$z_{k}:=\begin{cases}-\mathrm{Im}[\psi(1-i\lambda)]/\lambda&k=0\\ -\mathrm{Re}[\psi(1-i\lambda)]-\log h&k=1\\ \zeta(2-k)-\lambda^{2}z_{k-2}&k>1\end{cases},$ (17)
where $\psi$ is the digamma function, and where the parameter $\lambda$ is defined as
$\lambda:=\frac{d}{ch}.$
Remark 1.
The first series of (16) can be seen as a “singular error” that is a continuous extension from $d=0$ to $d>0$. The second series of (16) can be seen as a “jump error” that corresponds to the jump discontinuity between $d>0$ and $d<0$; this “jump error” does not appear when $d=0$ as the integral $I\left[\frac{g(x)}{c^{2}x^{2}}\right]$ is defined in the finite-part sense.
In order to prove Theorem 3, we will need the following two lemmas.
Lemma 4.
Define the family of rational zeta series
$f_{k}(z):=\sum_{m=0}^{\infty}z^{2m}\zeta_{h}(2m+2-k),\quad k=0,1,\dots$ (18)
where $\zeta_{h}$ is the modified zeta function (14). Then $f_{k}$ satisfies the recurrence relations
$f_{0}(z)$ $=\tfrac{\psi(1+z)-\psi(1-z)}{2z},$ (19)
$f_{1}(z)$ $=-\tfrac{\psi(1+z)+\psi(1-z)}{2}-\log h,$
$f_{k}(z)$ $=\zeta(2-k)+z^{2}f_{k-2}(z),$ $k>1.$
where $\psi$ is the digamma function. Note that (19) implies that each $f_{k}(z)$, as defined in (18), can be analytically continued to a meromorphic function on $\mathbb{C}$.
Proof.
The recurrence relation (19) is a direct consequence of the well-known expansion of the digamma function $\psi(1+z)$ as a rational zeta series [1, eq.(6.3.14)]. ∎
Lemma 5.
Using the notations (3-4), we have, for any integer $k\geq 0$,
$R_{h}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}-I_{R}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}=\begin{cases}\frac{1}{c^{2}}z_{k}h^{k-1}-(-1)^{k/2}\frac{d^{k-1}}{c^{k+1}}\frac{\pi}{2},&k\text{ even}\\ \frac{1}{c^{2}}z_{k}h^{k-1}+(-1)^{\frac{k-1}{2}}\frac{d^{k-1}}{c^{k+1}}\log d,&k\text{ odd}\end{cases}$ (20)
and
$L_{h}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}-I_{L}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}=\begin{cases}\frac{1}{c^{2}}z_{k}h^{k-1}-(-1)^{k/2}\frac{d^{k-1}}{c^{k+1}}\frac{\pi}{2},&k\text{ even}\\ -\frac{1}{c^{2}}z_{k}h^{k-1}-(-1)^{\frac{k-1}{2}}\frac{d^{k-1}}{c^{k+1}}\log d,&k\text{ odd}\end{cases},$ (21)
where $z_{k}$ is as defined in (17).
Proof.
Note that the series
$\frac{x^{k}}{d^{2}+c^{2}x^{2}}=\sum_{m=0}^{\infty}\frac{(-1)^{m}d^{2m}}{c^{2m+2}x^{2m+2-k}}$ (22)
converges whenever $\left|\frac{d}{cx}\right|<1$. To prove (20), we decompose the expression into a telescoping sum that consist of three components, $A_{k},B_{k}$, and $C_{k}$, as follows
$R_{h}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}$- $I_{R}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}=\underbrace{R_{h}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}-\sum_{m=0}^{\infty}\frac{(-1)^{m}d^{2m}}{c^{2m+2}}R_{h}\Big{[}\frac{1}{x^{2m+2-k}}\Big{]}}_{:=A_{k}}$ (23)
$\qquad+\underbrace{\sum_{m=0}^{\infty}\frac{(-1)^{m}d^{2m}}{c^{2m+2}}\left(R_{h}\Big{[}\frac{1}{x^{2m+2-k}}\Big{]}-f.p.I_{R}\Big{[}\frac{1}{x^{2m+2-k}}\Big{]}\right)}_{:=B_{k}}$
$\qquad+\underbrace{\sum_{m=0}^{\infty}\frac{(-1)^{m}d^{2m}}{c^{2m+2}}\left(f.p.I_{R}\Big{[}\frac{1}{x^{2m+2-k}}\Big{]}\right)-I_{R}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}}_{:=C_{k}}$
First, the series (22) implies that
$A_{k}=0,\quad k\geq 0$ (24)
for $\lambda=\frac{d}{ch}<1$. Next, by applying Corollary 2 to each term of the series in $B_{k}$, we have
$B_{k}=\sum_{m=0}^{\infty}\frac{(-1)^{m}d^{2m}}{c^{2m+2}}\zeta_{h}(2m+2-k)h^{-2m-1+k}=\frac{h^{k-1}}{c^{2}}\sum_{m=0}^{\infty}\left(i\lambda\right)^{2m}\zeta_{h}(2m+2-k)$ (25)
which is a rational zeta series. Then substituting $z=i\lambda$ in Lemma 4 implies that
$B_{k}=\frac{h^{k-1}}{c^{2}}z_{k}$ (26)
where $z_{k}$ is as defined in (17).
Finally, $C_{k}$ can be calculated by recursion as follows. Let
$C_{k}=\underbrace{\sum_{m=0}^{\infty}\frac{(-1)^{m}d^{2m}}{c^{2m+2}}\left(f.p.I_{R}\Big{[}\frac{1}{x^{2m+2-k}}\Big{]}\right)}_{:=C_{k}^{(1)}}-\underbrace{I_{R}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}}_{:=C_{k}^{(2)}}.$
Note that
$C_{k+2}^{(2)}$ $=\int_{0}^{a}\frac{x^{k+2}}{d^{2}+c^{2}x^{2}}\mathrm{d}x$
$=\frac{1}{c^{2}}\left(\int_{0}^{a}x^{k}\mathrm{d}x-d^{2}\int_{0}^{a}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\mathrm{d}x\right)$
$=\frac{1}{c^{2}}\left(\frac{a^{k+1}}{k+1}-d^{2}C_{k}^{(2)}\right)$
and similarly $C_{k+2}^{(1)}=\frac{1}{c^{2}}\left(\frac{a^{k+1}}{k+1}-d^{2}C_{k}^{(1)}\right)$. Thus $C_{k+2}=C_{k+2}^{(1)}-C_{k+2}^{(2)}=-\frac{d^{2}}{c^{2}}C_{k}$. The initial terms of $C_{k}$ are $C_{0}=-\frac{\pi}{2cd}$ and $C_{1}=\frac{\log d}{c^{2}}$ by straightforward calculations. Therefore we have
$C_{k}=\begin{cases}-(-1)^{k/2}\frac{d^{k-1}}{c^{k+1}}\frac{\pi}{2},&k\text{ even}\\ (-1)^{\frac{k-1}{2}}\frac{d^{k-1}}{c^{k+1}}\log d,&k\text{ odd}\end{cases}.$ (27)
Substituting (24,26,27) into (23) yields (20). Note that for fixed $h$, both sides of (20) are analytic functions of $\lambda=\frac{d}{ch}$, thus the identity also holds for $\lambda\geq 1$ by analytic continuation.
The proof of (21) is similar to that of (20). ∎
We are now ready to prove Theorem 3.
Proof of Theorem 3.
Using the Maclaurin series of $g(x)$, we have
$E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}x^{2}}\Big{]}=\sum_{k=0}^{\infty}\frac{g^{(k)}(0)}{k!}E_{h}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}.$ (28)
Each term of the above series can then be substituted by (20) and (21), which yields (16). ∎
Remark 2.
When the function $g(x)$ can be analytically continued to a neighborhood of $x=0$ in $\mathbb{C}$, the generalized Euler-Maclaurin formula (16) has a closed-form expression given by
$E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}x^{2}}\Big{]}=\frac{1}{c^{2}}\mathrm{Re}\Big{[}\frac{g(i\lambda h)-g(0)}{(i\lambda h)^{2}}\Big{]}h+\left(\frac{\pi}{cd}-\frac{2z_{0}}{c^{2}h}\right)\mathrm{Re}[g(i\lambda h)]$ (29)
For a derivation of (29), first of all, note that by the recurrence relation (17) the coefficients $z_{2k}$ in (16) satisfy $z_{2k}=(-\lambda^{2})^{k}(z_{0}+\frac{1}{2\lambda^{2}})$ when $k\geq 1$, thus
$\sum_{k=0}^{\infty}\frac{g^{(2k)}(0)}{(2k)!}h^{2k-1}\,z_{2k}$ $=\frac{g(0)z_{0}}{h}+\sum_{k=1}^{\infty}\frac{g^{(2k)}(0)}{(2k)!}h^{2k-1}\,(-\lambda^{2})^{k}(z_{0}+\tfrac{1}{2\lambda^{2}})$ (30)
$=-\frac{g(0)}{2\lambda^{2}h}+\left(\frac{z_{0}}{h}+\frac{1}{2\lambda^{2}h}\right)\sum_{k=0}^{\infty}\frac{g^{(2k)}(0)}{(2k)!}(i\lambda h)^{2k}$
$=-\frac{g(0)}{2\lambda^{2}h}+\left(\frac{z_{0}}{h}+\frac{1}{2\lambda^{2}h}\right)\mathrm{Re}[g(i\lambda h)]$
$=\frac{1}{2}\mathrm{Re}\Big{[}\frac{g(i\lambda h)-g(0)}{(i\lambda h)^{2}}\Big{]}h+\frac{z_{0}}{h}\mathrm{Re}[g(i\lambda h)].$
Second of all, the second series in (16) can be expressed as
$\sum_{k=0}^{\infty}(-\tfrac{d^{2}}{c^{2}})^{k}\frac{g^{(2k)}(0)}{(2k)!}=\mathrm{Re}[g(id/c)]=\mathrm{Re}[g(i\lambda h)].$ (31)
Combining (3031) yields (29). To see the connection of (29) with the existing generalized Euler-Maclaurin formulas in the literature, consider its limit as $d\to 0$ (thus $\lambda h=d/c\to 0$)
$\lim_{d\to 0}E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}x^{2}}\Big{]}=\frac{1}{c^{2}}\left(\frac{g^{\prime\prime}(0)}{2}h-\frac{2\zeta(2)}{h}g(0)\right)+\lim_{d\to 0}\frac{\pi}{cd}\,[g(0)+O(d^{2})]$ (32)
This agrees with the finite-part quadrature correction of [19, eq.(3.2)-(3.4)] :
$E_{h}\Big{[}\frac{g(x)}{x^{2}}\Big{]}=\frac{g^{\prime\prime}(0)}{2}h-\frac{2\zeta(2)}{h}g(0)=\left(-\frac{2\zeta(2)}{h^{2}}\frac{g(0)}{0!}-\frac{2\zeta(0)}{h^{0}}\frac{g^{\prime\prime}(0)}{2!}\right)h$ (33)
and the fact that there is a jump $\frac{\pi}{d}\,g(0)$ between the classical and principal value integrals.
3.2 Euler-Maclaurin formula for an off-mesh near singularity
The generalized Euler-Maclaurin formula of Theorem 3 can be further extended to an integral $I\big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\big{]}$ where $x_{s}\neq 0$, whose near-singularity at $x_{s}$ is no long a mesh point of the Trapezoidal rule $T_{h}$; however, without loss of generality, we assume $x_{s}$ is small enough such that $x=0$ is still the closest mesh point from the near-singularity.
Theorem 6 (Off-mesh Euler-Maclaurin for near-singular integral).
Given $g(x)\in C^{\infty}([a,a])$ and $c,d\neq 0$,
$E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}\sim-\frac{1}{c^{2}}\sum_{k=0}^{\infty}\frac{g^{(k)}(x_{s})}{k!}h^{k-1}p_{k,s}+\frac{\pi}{cd}\sum_{k=0}^{\infty}\frac{g^{(2k)}(x_{s})}{(2k)!}\left(-\frac{d^{2}}{c^{2}}\right)^{k}$ (34)
as $h\to 0$ and $d\to 0$, where the numbers $p_{k,s}$ satisfy the recurrence relation
$p_{k,s}=\begin{cases}-\mathrm{Im}[\psi(1-s-i\lambda)+\psi(1+s-i\lambda)]/\lambda&k=0\\ -\mathrm{Re}[\psi(1-s-i\lambda)-\psi(1+s-i\lambda)]&k=1\\ -(-s)^{k-2}-\lambda^{2}p_{k-2,s}&k\geq 2\end{cases}$ (35)
where $\psi$ is the digamma function, and where the parameter $\lambda=d/(ch)$.
Note that $\lim_{s\to 0}p_{k,s}=z_{k}+(-1)^{k}z_{k}$, so the formula (34) reduces to (16) as $x_{s}\to 0$. In order to prove Theorem 6, we first need the next three lemmas.
Lemma 7.
Using the notations (3) and (7), we have, for any integer $k\geq 0$,
$R_{h,s}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}-I_{R}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}=\begin{cases}\frac{1}{c^{2}}z_{k,s}h^{k-1}-(-1)^{k/2}\frac{d^{k-1}}{c^{k+1}}\frac{\pi}{2},&k\text{ even}\\ \frac{1}{c^{2}}z_{k,s}h^{k-1}+(-1)^{\frac{k-1}{2}}\frac{d^{k-1}}{c^{k+1}}\log d,&k\text{ odd}\end{cases}$ (36)
and
$L_{h,s}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}-I_{L}\Big{[}\frac{x^{k}}{d^{2}+c^{2}x^{2}}\Big{]}=\begin{cases}\frac{1}{c^{2}}z_{k,-s}h^{k-1}-(-1)^{k/2}\frac{d^{k-1}}{c^{k+1}}\frac{\pi}{2},&k\text{\,even}\\ -\frac{1}{c^{2}}z_{k,-s}h^{k-1}-(-1)^{\frac{k-1}{2}}\frac{d^{k-1}}{c^{k+1}}\log d,&k\text{\,odd}\end{cases}$ (37)
where the coefficients $z_{k,s}$ satisfy the recurrence relation
$z_{k,s}=\begin{cases}-\mathrm{Im}\Big{[}\psi(1+s-i\lambda)\Big{]}/\lambda&k=0\\ -\mathrm{Re}[\psi(1+s-i\lambda)]-\log h&k=1\\ \zeta(2-k,1+s)-\lambda^{2}\,z_{k-2,s}&k>1\end{cases}$ (38)
Proof.
The proof of Lemma 7 follows similar arguments as in the proof of Lemma 5. In particular, based on Theorem 1 a series expression for $z_{k,s}$ similar to (25) can be obtained, giving
$z_{k,s}=\sum_{m=0}^{\infty}\left(i\lambda\right)^{2m}\zeta_{h}(2m+2-k,1+s).$ (39)
Then one can show that $z_{k,s}$ defined by the series (39) satisfies the recurrence relation (38) by substituting $z=i\lambda$ in Lemma 8 below. ∎
Lemma 8.
Define the family of rational zeta series
$f_{k,s}(z):=\sum_{m=0}^{\infty}z^{2m}\zeta_{h}(2m+2-k,1+s),\quad k=0,1,\dots$ (40)
where $\zeta_{h}(z,s)$ is the modified Hurwitz zeta function (10). Then $f_{k}$ satisfies the recurrence relation
$f_{0,s}(z)$ $=\tfrac{\psi(1+s+z)-\psi(1+s-z)}{2z},$ (41)
$f_{1,s}(z)$ $=-\tfrac{\psi(1+s+z)+\psi(1+s-z)}{2}-\log h,$
$f_{k,s}(z)$ $=\zeta(2-k,1+s)+z^{2}f_{k-2,s}(z)$ $\text{for }k>1.$
where $\psi$ is the digamma function. Note that (41) implies that each $f_{k,s}(z)$, as defined in (40), can be analytically continued to a meromorphic function on $\mathbb{C}$.
Proof.
The connection between the rational zeta series (40) and the digamma function $\psi$ can be established by the relation [16, eq.(25.11.12)]
$\zeta(n+1,a)=\frac{(-1)^{n+1}\psi^{(n)}(a)}{n!},\quad n=1,2,3,\dots$ (42)
Lemma 9.
For any integer $k\geq 0$,
$\zeta(-k,1+s)+(-1)^{k}\zeta(-k,1-s)=-s^{k}$ (43)
Proof.
Let $B_{n}(x)$ be the Bernoulli polynomial of degree $n$. By the identities [1]
$\zeta(-n,x)$ $=-\frac{B_{n+1}(x)}{n+1},$
$B_{n}(1+x)$ $=B_{n}(x)+nx^{n-1}$
$B_{n}(1-x)$ $=(-1)^{n}B_{n}(x),$
we have
$\zeta(-k,1+s)+(-1)^{k}\zeta(-k,1-s)$ $=-\frac{1}{k+1}\Big{(}B_{k+1}(1+s)+(-1)^{k}B_{k+1}(1-s)\Big{)}$
$=-\frac{1}{k+1}\Big{(}\big{(}B_{k+1}(s)+(k+1)s^{k}\big{)}-B_{k+1}(s)\Big{)}$
$=-s^{k}$
We are now ready to prove Theorem 6.
Proof of Theorem 6.
If we define $\hat{x}:=x-x_{s}$, then
$E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}=E_{h,-s}\Big{[}\frac{g(\hat{x}+x_{s})}{d^{2}+c^{2}\hat{x}^{2}}\Big{]}=\sum_{k=0}^{\infty}\frac{g^{(k)}(x_{s})}{k!}E_{h,-s}\Big{[}\frac{\hat{x}^{k}}{d^{2}+c^{2}\hat{x}^{2}}\Big{]}$ (44)
then substituting (3637) into the above series yields
$E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}=-\frac{1}{c^{2}}\sum_{k=0}^{\infty}\frac{g^{(k)}(x_{s})}{k!}h^{k-1}(z_{k,-s}+(-1)^{k}z_{k,s})$ (45)
$+\frac{\pi}{cd}\sum_{k=0}^{\infty}\frac{g^{(2k)}(x_{s})}{(2k)!}\Big{(}-\frac{d^{2}}{c^{2}}\Big{)}^{k}.$
Then using the recurrence (38) and Lemma 9, it immediately follows that the coefficients
$z_{k,-s}+(-1)^{k}z_{k,s}\equiv p_{k,s}$ (46)
satisfy the recurrence relation (35). ∎
Remark 3.
Similar to (29), when $g(x)$ is analytic in a neighborhood of $x=x_{s}$ in $\mathbb{C}$ the generalized Euler-Maclaurin formula (34) has the closed form expression
$E_{h}\Big{[}\frac{g(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\Big{]}$ $=-\frac{1}{c^{2}h}\Bigg{\{}\left(p_{0,s}+\frac{1}{s^{2}+\lambda^{2}}\right)\mathrm{Re}[g(x_{s}+i\lambda h)]-\frac{g(x_{s}-sh)}{\lambda^{2}+s^{2}}$ (47)
$\quad+\left(p_{1,s}-\frac{s}{s^{2}+\lambda^{2}}\right)\frac{\mathrm{Im}[g(x_{s}+i\lambda h)]}{\lambda}\Bigg{\}}+\frac{\pi}{cd}\mathrm{Re}[g(x_{s}+i\lambda h)].$
This expression can be derived as follows.
  1. 1.
    The recurrence relation (35) implies that $p_{k,s}$ has the following closed-form expression
    $p_{2k,s}$ $=-\frac{s^{2k}-(-\lambda^{2})^{k}}{s^{2}+\lambda^{2}}+(-\lambda^{2})^{k}p_{0,s},$ (48)
    $p_{2k+1,s}$ $=\frac{s^{2k+1}-(-\lambda^{2})^{k}s}{s^{2}+\lambda^{2}}+(-\lambda^{2})^{k}p_{1,s}.$
    Substituting (48) into the first series on the right-hand side of (34), we have
    $\sum_{k=0}^{\infty}$ $\frac{g^{(k)}(x_{s})}{k!}h^{k-1}p_{k,s}$ (49)
    $=\sum_{k=0}^{\infty}\frac{g^{(2k)}(x_{s})}{(2k)!}h^{2k-1}\left(\frac{(-\lambda^{2})^{k}-s^{2k}}{s^{2}+\lambda^{2}}+(-\lambda^{2})^{k}p_{0,s}\right)$
    $\quad+\sum_{k=0}^{\infty}\frac{g^{(2k+1)}(x_{s})}{(2k+1)!}h^{2k}\left(\frac{s^{2k+1}-(-\lambda^{2})^{k}s}{s^{2}+\lambda^{2}}+(-\lambda^{2})^{k}p_{1,s}\right)$
    $=\frac{1}{(s^{2}+\lambda^{2})h}\left(\mathrm{Re}[g(x_{s}+i\lambda h)]-g(x_{s}-s\,h)-s\frac{\mathrm{Im}[g(x_{s}+i\lambda h)]}{\lambda}\right)$
    $\quad+\frac{p_{0,s}}{h}\mathrm{Re}[g(x_{s}+i\lambda h)]+\frac{p_{1,s}}{h}\frac{\mathrm{Im}[g(x_{s}+i\lambda h)]}{\lambda}.$
  2. 2.
    The second series on the right-hand side of (34) can be expressed as
    $\sum_{k=0}^{\infty}(-\tfrac{d^{2}}{c^{2}})^{k}\frac{g^{(2k)}(x_{s})}{(2k)!}=\mathrm{Re}[g(x_{s}+id/c)]=\mathrm{Re}[g(x_{s}+i\lambda h)].$ (50)
    Now, combining the above expressions, (34) becomes (47).
3.3 New singular Trapezoidal rule for off-mesh singularity
As a consequence of Theorem 6, we obtain in the following corollary an off-mesh version of the generalized Euler-Maclaurin formula for a finite-part (hypersingular) integral, which seems to be new.
Corollary 10 (Off-mesh hypersingular Euler-Maclaurin formula).
$E_{h}\Big{[}\frac{g(x)}{(x-x_{s})^{2}}\Big{]}$ $=\frac{g(0)-g(x_{s})-g^{\prime}(x_{s})(-x_{s})}{x_{s}^{2}}h$ (51)
$\quad-\frac{\zeta(2,1-s)+\zeta(2,1+s)}{h}g(x_{s})$
$\quad-(-\psi(1-s)+\psi(1+s))g^{\prime}(x_{s})$
Proof.
Using the definition (35) and the relation (42), we have
$\lim_{\lambda\to 0}p_{0,s}=\zeta(2,1-s)+\zeta(2,1+s)$ (52)
$\lim_{\lambda\to 0}p_{1,s}=-\psi(1-s)+\psi(1+s)$
$\lim_{\lambda\to 0}\mathrm{Re}\Big{[}\tfrac{g(x_{s}+i\lambda h)}{i\lambda h}\Big{]}=g^{\prime}(x_{s})$
Substituting the above limits into (47) as $d\to 0$ (hence $\lambda\to 0$) yields (51), where the jump $\lim_{d\to 0}\frac{\pi}{cd}\mathrm{Re}[g(x_{s}+i\lambda h)]$ is ignored to be consistent with the value of the finite-part integral. ∎
Remark 4.
The following quadrature for a finite-part integral is presented in [19, eq.(3.10)]
$I\Big{[}\frac{g(x)}{x^{2}}\Big{]}\approx\tilde{T}_{h,1/2}\Big{[}\frac{g(x)}{x^{2}}\Big{]}-\pi^{2}h^{-1}g(h/2),$ (53)
which is spectrally accurate; here, $\tilde{T}$ represents an ordinary Trapezoidal rule (not punctured). One can derive this formula as a consequence of Corollary 10. Let $s\to 1/2$ (hence $x_{s}\to h/2$) in (51) and using the fact that [16, eq.(25.11.8), eq.(5.15.5)]
$\lim_{s\to\frac{1}{2}}\zeta(2,1-s)+\zeta(2,1+s)=\pi^{2}-4$ (54)
$\lim_{s\to\frac{1}{2}}-\psi(1-s)+\psi(1+s)=2$
we arrive at the expression
$E_{h}\Big{[}\frac{g(x)}{(x-h/2)^{2}}\Big{]}=\frac{g(0)}{(h/2)^{2}}h-\pi^{2}h^{-1}g(h/2)$ (55)
If the $g(0)$ term in (55) is included back into the punctured Trapezoidal rule on the left-hand side to yield the ordinary Trapezoidal rule, then the error expression becomes
$\tilde{E}_{h}\Big{[}\frac{g(x)}{(x-h/2)^{2}}\Big{]}\equiv\tilde{E}_{h,1/2}\Big{[}\frac{g(x)}{x^{2}}\Big{]}=-\pi^{2}h^{-1}g(h/2),$ (56)
where $\tilde{E}$ denotes the error for $\tilde{T}$ in (53), which is the consistency we wanted to show.
4 Numerical Examples
We present convergence results for the generalized Euler-Maclaurin formulas of Section 3.
Let $\mathrm{Ei}(z)$ be the exponential integral and consider the near-singular integral
$I=\int_{-1}^{1}\frac{de^{x}}{d^{2}+x^{2}}\,\mathrm{d}x$ (57)
whose exact value
$I=\mathrm{Im}\left\{e^{id}\big{(}\mathrm{Ei}(1-id)-\mathrm{Ei}(-1-id)\big{)}\right\}$
is well-defined for all $d\in\mathbb{R}$. Note that the integral has a well-defined limit $I\to\pm\pi$ as $d\to 0^{\pm}$. Figure 2 presents convergence results of the corrected trapezoidal rule based on (16), with $h=2/N$, applied to the integral (57) for different values of $d$, where the near-singular point $x=0$ is a quadrature node; the correction is based on the closed-form expression (29), so the corrected quadrature converges exponentially. Machine precision is reached around $N=100$ independent of the size of $d$. On the contrary, the convergence of the regular Trapezoidal rule without correction slows down significantly as $d\to 0$ and the integral becomes closer to singular.
Figure 2: Convergence of the Trapezoidal rule (mesh size $h=2/N$), with and without corrections, applied to the integral of $f(x)=\frac{de^{x}}{d^{2}+x^{2}}$, for $d=0.1,0.01,0.0001$. The near-singular correction is based on the formula (29).
Similar convergence results are observed for the off-mesh near-singularity in Figure 3, where the corrected trapezoidal rule based on (34) and the closed-form correction formula (47) is applied to the integral
$I=\int_{-1}^{1}\frac{de^{x}}{d^{2}+c^{2}(x-x_{s})^{2}}\,\mathrm{d}x=\mathrm{Im}\left\{e^{x_{s}+\frac{id}{c}}\big{(}\mathrm{Ei}(1-x_{s}-\tfrac{id}{c})-\mathrm{Ei}(-1-x_{s}-\tfrac{id}{c})\big{)}\right\}$
with $x_{s}=0.1$. Note that $x_{s}$ is not a quadrature node in this example.
Figure 3: Similar to Figure 2 but for the integral of $f(x)=\frac{de^{x}}{d^{2}+c^{2}(x-x_{s})^{2}}$ with near-singularity at $x_{s}=0.1$ (not a quadrature node). The correction is based on the formula (47).
Remark 5.
For a general integral where the analytic expression of $g(x)$ is unknown or where $g(x)$ cannot be analytically continued to a neighborhood of the near-singular point, a truncation of the error corrections (16) and (34) can be computed by approximating $g^{(k)}(0)$ or $g^{(k)}(x_{s})$ using finite difference formulas on a small stencil centered at $x=0$. In practice, we find that convergence results that are almost identical to Figures 2 and 3 can be achieved by including correction terms in (16) and (34) that involve up to the $6^{\rm th}$ derivative of $g(x)$. Matlab code for these computations are available at https://github.com/bobbielf2/PolygammaTrap.
5 Conclusion
We have presented generalized Euler-Maclaurin formulas for near-singular integrals of the form
$\int\frac{f(x)}{d^{2}+c^{2}(x-x_{s})^{2}}\,\mathrm{d}x,$
which achieve high-order convergence for any values of $d$.
A future direction is to further extend the current work to integrals of the form
$\int\frac{f(x)}{(d^{2}+c^{2}(x-x_{s})^{2})^{p}}\,\mathrm{d}x,~~~p>1$
and develop near-singular quadrature rules for integral operators such as (2), which are near-singular integrals on curves in $\mathbb{R}^{2}$ that arise in the solution of partial differential equations. Another future direction is to derive analogous generalizations of the Euler-Maclaurin formula in $\mathbb{R}^{2}$ and apply them to surface integrals in $\mathbb{R}^{3}$, which will be an extension of the singular quadrature method of [21, 23].